Maria is 5 years older than ralph . Richard is twice as old as ralph . The sum oftheir ages is 65 how old is each?
Answer by algebra hello(55)

let Ralph's age=x Maria's age=x+5 Richard's age=2x according to question,x+2x+x+5=65 3x=65-5 3x=60 x=20